A bank contains 300 coins, all of them quarters or dimes. If the money totals $67.50, how many of each coin are in the bank?
Answer by chriswen(106)

Let x be the number of quarters.
Let 300-x be the number of dimes.
...
25x+10(300-x)=6750
25x+3000-10x=6750
15x=6750-3000
15x=3750
x=250
300-x=50
...
Therefore, there are 250 quarters and 50 dimes in the bank.